Jonathan borrowed $10,000 at 6% annual
compound interest. He agreed to repay the loan with
five equal annual payments at end-of-years 1–5. How
much of the annual payment is interest, and how much
principal is there in each annual payment?
1
Expert's answer
2020-08-04T16:06:50-0400
"\frac {1}{6} \times 10000 \times (1+0.1)^6=2952.6"